Assume that the resources best sulted to producing a partícular service are preferentially used in the production of that service and that as the economy moves down along the production possibiNties frontlier, one worker at a time is transferred from mowing lawns to washing cars. Using the blue points (circle symbol), graph the production possibilities frontier (PPF) for this economy on the following graph. Then use the black point (plus symbol to identify point A, the greeภ point triangle Symbol to identify point B the orange point square symbol) to identify pont C and the on point (diamond symbol) to identily point D on the graph. 50 PPF 35 30 25 20 าธ 10 A-Z 0 5 10 20 25 30 35 40 45 50 Quantity of Cars Washed True or False: The production possibilities frontier consists of straight-line segments, rather than being smoothly bowed outward throughout, because each worker faces a constant trade-off between mowing lans and washing cars. TrueTrue or False: The production possibilities frontier consists of straight-line segments, rather than being smoothly bowed outward throughout, because each worker faces a constant trade-off between mowing lawns and washing cars. True False Indicate whether each of the following allocations is efficient or inefficient. Allocation Efficient Inefficient A-Z Grade It Now Save & Continue Continue without saving
Answer #1

(A) Lawn mowed by Andrew, Darnell and Jacques = 2 x 10 hours + 1 x 10 + 1 x 10 = 20 + 10 + 10 = 40

(B) Car washed by all = 1 x 10 + 1 x 10 + 2 x 10 = 40

(C) Lawn mowed = 2 x 5 hours + 1 x 5 + 1 x 5 = 10 + 5 + 5 = 20

Car washed = 1 x 5 hours + 1 x 5 + 2 x 5 = 20

(D) Lawn mowed = 2 x 5 hours + 1 x 10 hours = 20

Car washed = 1 x 5 + 1 x 10 = 15

Opportunity cost is the value of next best alternative foregone.

1. Andrew:

1 car wash = 2 lawn mowed

It implies opportunity cost of washing car is 2 Lawn mowed.

2. Darnell:

1 car wash = 1 lawn mowed

Opportunity cost of 1 car wash is 1 lawn mowed.

3. Jacques:

2 car wash = 1 lawn mowed

1 car wash = 1/2 = 0.5 lawn mowed

Opportunity cost is 0.5 lawn mowed.

50 T PPF Σ 30 e 20 10 0 5 10 20 25 30 35 40 45 50 Quantity af Cars Washed

True.

A, B, C - These are efficient because lie on the PPF.

D is inefficient because it lies below PPF which means some resources are left idle.